# The 1/3 Geometric Constant: Scale Invariance and the Origin of ’Missing Energy’ in 3D Quantum Fragmentation

### v4

@@ Abstract
We report the discovery of a universal geometric constraint on the detection of kinetic energy release (KER) in three-dimensional quantum fragmentation. By analyzing the dissociation of localized wavepackets, we demonstrate that the $4\pi r^{2}$ radial volume element acts as a topological filter that inherently masks a significant portion of a system’s energy budget, imposing a fundamental peak-to-mean bound of $R_{E}<0.5$. We introduce an invariant scaling law, $\alpha=MQ/\zeta$, and prove that the resulting energy detection ratio is scale-invariant across twelve orders of magnitude, bridging attosecond molecular science and nuclear physics. We identify a universal “Geometric Zero-Point” at $R_{E}\approx 0.33$, which precisely replicates the 7 eV discrepancy in $H_{2}^{+}$ fragmentation. Furthermore, we show that the population of excited-state manifolds and the increase in nuclear localization ($\zeta$) provide a definitive geometric mechanism for the spectral broadening observed across atomic and subatomic scales. Remarkably, the spectral morphology derived from our scaling law aligns with the universal 1/3 energy landmark of historical beta decay, while the high-mass limit naturally accounts for the sharpening of alpha spectra. Our results suggest that “missing energy” is often a topological artifact of 3D geometry rather than an exclusive signature of undetected particles. This work establishes a universal master curve for energy reconstruction and identifies a “detection crisis” in highly localized systems, where the true interaction energy becomes effectively invisible to peak-centric calorimetry.
@@ I Introduction
A central challenge in experimental physics is the reconstruction of the total energy budget from a detected energy distribution. Across disparate scales—from the electron-volts of molecular dissociation to the mega-electron-volts of nuclear transitions—a persistent anomaly remains: the primary spectral peak often captures only a fraction of the available energy. Historically, this “missing energy” has been the catalyst for some of the most profound discoveries in physics, most notably the hypothesis of the neutrino to resolve the energy deficit in beta decay Ellis and Wooster (1927).
As illustrated in our conceptual framework (Fig. 1), we propose that these discrepancies are linked by a universal topological mechanism. The Geometric Masking effect (Fig. 1, Row 1) arises from the inherent $4\pi r^{2}$ volume weighting of 3D space, which suppresses detection probability near the coordinate origin where potential energy is maximal. This results in a systematic shift of the detected kinetic energy release (KER) peak away from the true system mean.
In this work, we demonstrate that these anomalies share a common geometric origin defined by an invariant scaling law, $\alpha=MQ/\zeta$ (Fig. 1, Row 2). We prove that for any localized 3D wavefunction, the interplay between mass ($M$), localization ($\zeta$), and repulsive charge ($Q$) constrains the energy detection ratio. We identify a 0.33 geometric landmark—a universal attractor where the detected peak represents approximately one-third of the total energy budget. This scaling law predicts that as particles become more localized, the detected energy decays precipitously, creating a “detection crisis” where the true interaction energy becomes effectively invisible to peak-centric calorimetry.
The predictive power of this framework is underscored by the striking alignment between our scaling law and diverse experimental anomalies (Fig. 1, Row 3). In attosecond calorimetry, our model accurately predicts the observed 7 eV “missing energy” in $H_{2}^{+}$ fragmentation, yielding a detection ratio of $R_{E}\approx 0.32$ that matches modern experimental data Zhu (2020, 2021) and accounts for the spectral broadening induced by excited-state populations. This molecular landmark is mirrored almost perfectly in the subatomic regime; our $P(E)$ distribution provides a robust fit for the historical 210Bi beta decay spectrum within the non-relativistic limit, where the average electron energy remains tethered to approximately one-third of the endpoint energy Ellis and Wooster (1927). The persistence of this 0.33 ratio across twelve orders of magnitude, regardless of the specific interaction or mass scale, identifies the geometric volume element as the universal arbiter of energy detection. Ultimately, this framework provides a unified baseline for energy reconstruction, offering a robust geometric filter to distinguish between the intrinsic masking effects of 3D quantum geometry and the true signatures of new physics.
@@
II The Geometric Physics of Fragmentation
@@
The core of our model rests on the sudden approximation, which remains valid when a system’s Hamiltonian changes on a timescale $\tau_{perturb}\ll\hbar/\Delta E$. Under this condition, the initial wavefunction $\psi_{0}(\mathbf{r})$ remains “frozen” at $t=0$, as there is insufficient time for the probability density to reorganize Makarov (2016). In the context of our investigation, we model the fragmentation as an instantaneous transition from an attractive potential to a repulsive Coulomb configuration:
$V_{initial}(r)=-\frac{Z_{att}}{r}\longrightarrow V_{final}(r)=+\frac{Q}{r}.$ (1)
@@
This approximation captures the essential dynamics of ultrafast processes across scales, from attosecond molecular dissociation to the loss of confinement in nuclear physics. To bridge the gap between initial state geometry and final energy observables, we utilize the de Broglie-Bohm formulation. The total energy $E(\mathbf{r})$ of a fragment at a specific coordinate $\mathbf{r}$ is the sum of the repulsive potential and the local kinetic energy density $T_{local}(\mathbf{r})$. For a Slater-type orbital $R(r)=Nr^{n-1}e^{-\zeta r}$, the analytical expression for the conserved total energy is:
$E(r)=\frac{1}{2M}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)+\frac{Q}{r}.$ (2)
@@
The observable energy distribution $P(E)$ is constructed by sampling initial positions according to the 3D radial probability density $\rho(r)=4\pi r^{2}|\psi_{0}(r)|^{2}$. Crucially, this $r^{2}$ volume weighting acts as a topological filter. In highly localized systems (high $\zeta$), the majority of the probability density is concentrated at small $r$, where the local kinetic energy is highly negative. This results in an energy distribution where the most probable detected energy ($E_{peak}$) is shifted significantly lower than the mean system energy.
III Numerical Evolution and Parametric Sensitivity
@@
To evaluate the predictive power of our 3D Bohmian-based model, we calculated the resulting energy distributions $P(E)$ and compared them against full quantum simulations. While detailed comparisons with tRecX Scrinzi (2022) Time Dependent Schrödinger Equation (TDSE) benchmarks are provided in Appendix A.2, we note that our analytical local-energy approach precisely captures the primary experimental signatures—specifically the peak energy $E_{\text{peak}}$—across varying regimes. Notably, the consistency in peak prediction (see Fig. 5(a)) confirms that our simplified method effectively captures the essential physics of the 3D Coulomb explosion.
III.1 Parametric Sensitivity and Distribution Morphology
@@
To characterize the influence of physical parameters on the detected energy, we systematically varied the repulsive Coulomb charge $Q$ and the orbital exponent $\zeta$ (see Fig. 2). As illustrated in Fig. 2(a), increasing the repulsive charge $Q$ leads to a linear shift in the distribution $P(E)$ toward higher energies. This is accompanied by a moderate broadening of the spectral width, as a larger $Q$ increases the potential gradient across the spatial extent of the orbital.
In Fig. 2(b), we examine the effect of the orbital exponent $\zeta$. As $\zeta$ increases, the wavefunction becomes more tightly bound, simultaneously increasing the local kinetic energy density ($T_{\text{local}}\propto\zeta^{2}$) and placing the particle density in regions of higher potential energy. Interestingly, the distribution broadens much more rapidly with respect to $\zeta$ than with $Q$. This suggests that the internal “quantum pressure” of the initial state plays a more dominant role in determining the final energy dispersion than the external repulsive force itself. The shape of these curves confirms that fragment detectors fundamentally probe the radial geometry of the initial state rather than the integrated energy mean.
III.2 The Energy Detection Ratio $R_{E}$ and the Detection Crisis
To quantify the discrepancy between experimentally observed fragment signals and theoretical energy conservation, we define the dimensionless energy detection ratio:
@@ III.2 The Energy Detection Ratio $R_{E}$ and the Detection Crisis
where $E_{\text{peak}}$ is the mode of the distribution $P(E)$ and $\langle E\rangle$ is the integrated expectation value of the total energy. This ratio serves as a metric for the efficiency with which a detector captures the system’s total energy budget. Fig. 2(c) and 2(d) illustrates the systematic evolution of $R_{E}$ across the parametric space.
III.2.1 High-Charge Convergence and Classicalization
As shown by the black curve in Fig. 2(c), $R_{E}$ increases monotonically with the repulsive charge $Q$, asymptotically approaching a limit of $0.5$. In the high-$Q$ regime, the repulsive potential $V=Q/r$ dominates the local kinetic energy terms, “classicalizing” the landscape and shifting the peak of the distribution closer to the statistical mean. However, the fact that $R_{E}$ remains strictly bounded below $0.5$ underscores that the 3D quantum spatial spread is an inescapable constraint on the fragment energy spectrum.
III.2.2 The Shell-Dependent Detection Crisis
A precipitous decline in $R_{E}$ is observed as the orbital exponent $\zeta$ increases (red curve in Fig. 2(c)). In our framework, $\zeta$ serves as a proxy for the effective nuclear charge and localization; small $\zeta$ values represent diffuse valence orbitals, while large $\zeta$ characterize highly localized core orbitals. As $\zeta$ increases, the quadratic scaling of the “quantum pressure” near the nucleus ($T_{\text{local}}\propto\zeta^{2}$) drives the integrated mean energy $\langle E\rangle$ to values much higher than the peak $E_{\text{peak}}$, which is primarily determined by the most probable radial position.
Our findings reveal a “shell-dependent detection crisis.” For valence-shell Coulomb explosions (low $\zeta$), $R_{E}$ is relatively high, indicating that the detected peak provides a reasonably transparent window into the system’s potential energy. Conversely, inner-shell events (high $\zeta$) are subject to a massive geometric shift; for $\zeta\approx 5$, the detector may observe as little as 15% of the total energy mean. Given that localization typically scales with particle mass, the extreme confinement of nucleons in nuclear physics implies a significantly suppressed $R_{E}$. The global contour map in Fig. 2(d) synthesizes these trends, confirming that the deeper a particle is “buried” in the atomic potential, the more the 3D quantum geometry masks the true interaction energy from classical detection.
IV The Dimensionless Master Curve and Scaling Symmetry
@@
The emergence of the 0.33 landmark is rooted in a fundamental scaling symmetry of three-dimensional fragmentation dynamics. To characterize this, we examine the peak-to-mean ratio $R_{E}$ across the parametric landscape of mass ($M$), orbital localization ($\zeta$), and repulsive charge ($Q$). As demonstrated in Fig. 3, the ratio is governed by a single dimensionless coupling constant $\alpha$ defined by:
@@
For a $1s$ Slater orbital, substituting the scaling relation for the dimensionless coupling reveals that the local kinetic density scales as $\rho_{T}\propto\zeta^{2}/M=\frac{1}{\alpha}\zeta Q$, while the local potential density follows $\rho_{V}\propto Q\zeta$ (see Sec. A.4). By transforming to a dimensionless radial grid $r_{s}=\zeta r$, both energy components at every coordinate point scale linearly with the product $\zeta Q$ for any fixed $\alpha$. This local synchronization ensures that the geometric morphology of the resulting $P(E)$ distribution is invariant to the absolute energy scale, depending solely on the coupling constant $\alpha$. This is confirmed in Fig. 3(a), where normalized distributions $P(E/Q\zeta)$ for varying $Q$ and $\zeta$ collapse onto identical master curves. As $\alpha\to\infty$, the distribution narrows toward a fixed peak position, providing a robust, scale-invariant baseline for extracting intrinsic parameters ($M,\zeta,Q$) from fragmented experimental spectra.
Furthermore, this symmetry demonstrates that the detection ratio $R_{E}$ remains invariant along $M\propto\zeta$ trajectories. The system undergoes a self-similar expansion in energy space where the $4\pi r^{2}$ masking effect is preserved as a constant geometric partition. The invariance observed in the parallel trajectories of Fig. 3(b) and the invariant planes of Fig. 3(c) is therefore a direct consequence of this homogeneous scaling. At the $\alpha\approx 1$ equilibrium, this collapse manifests as a horizontal 1D plateau at $R_{E}\approx 0.33$, identifying the landmark as a structural constant of 3D quantum geometry rather than a dynamical variable.
@@
As shown in the master curve in Fig. 3(d), the numerical behavior of $R_{E}$ is characterized by the following limits:
@@
When $\alpha=1$, the system achieves a “geometric equilibrium” where the $4\pi r^{2}$ volume element sequesters exactly two-thirds of the total energy in the high-momentum radial tail, yielding the 0.33 landmark. As $\alpha$ increases (representing high $Q$ or low $M$), the ratio monotonically climbs toward the 0.5 universal bound, a regime where the repulsive field dominates the initial-state quantum pressure.
@@
Conversely, as $\alpha$ increases (representing high localization $\zeta$), $R_{E}$ decays precipitously, falling below 0.2 for $\zeta>5$. This signals a profound “detection crisis” where the vast majority of interaction energy becomes experimentally invisible to peak-centric calorimetry. This confirms that for sufficiently localized states—such as core-shell electrons or compressed nucleons—the true interaction energy is effectively masked by the topological constraints of 3D space. Consequently, the mapping $R_{E}(\alpha)$ serves as a universal master curve for energy reconstruction across all scales of 3D quantum fragmentation.
V Discussion: From Molecular Calorimetry to Beta Decay
V.1 Empirical Validation: The $H_{2}^{+}$ Energy Discrepancy

### v5

@@ Abstract
We report the discovery of a universal geometric constraint on the detection of kinetic energy release (KER) in three-dimensional quantum fragmentation. By analyzing the dissociation of localized wavepackets, we demonstrate that the $4\pi r^{2}$ radial volume element acts as a topological filter that inherently masks a significant portion of a system’s energy budget, imposing a fundamental peak-to-mean bound of $R_{E}<0.5$. We introduce an invariant scaling law, $\alpha=MQ/\zeta$, and prove that the resulting energy detection ratio is scale-invariant across twelve orders of magnitude, bridging attosecond molecular science and nuclear physics. We identify a universal geometric landmark at $R_{E}\approx 0.33$, which precisely replicates the 7 eV discrepancy in $H_{2}^{+}$ fragmentation. Furthermore, we show that the population of excited-state manifolds and the increase in nuclear localization ($\zeta$) provide a definitive geometric mechanism for the spectral broadening observed across atomic and subatomic scales. Remarkably, the spectral morphology derived from our scaling law aligns with the universal 1/3 energy landmark of historical beta decay, while the high-mass limit naturally accounts for the sharpening of alpha spectra. Our results suggest that “missing energy” is often a topological artifact of 3D geometry rather than an exclusive signature of undetected particles. This work establishes a universal master curve for energy reconstruction and identifies a “detection crisis” in highly localized systems, where the true interaction energy becomes effectively invisible to peak-centric calorimetry.
@@ I Introduction
A central challenge in experimental physics is the reconstruction of a system’s total energy budget from detected spectral distributions. Across twelve orders of magnitude—from the electron-volts of molecular dissociation to the mega-electron-volts of nuclear transitions—a persistent anomaly remains: the primary spectral peak often captures only a fraction of the available interaction energy. Historically, such “missing energy” has served as a catalyst for fundamental discoveries, most notably the neutrino hypothesis to resolve the energy deficit in $\beta$ decay Ellis and Wooster (1927).
@@
In this Article, we propose that these disparate anomalies share a common topological origin. As illustrated in our conceptual framework (Fig. 1), we identify a Geometric Masking effect arising from the inherent $4\pi r^{2}$ volume weighting of three-dimensional space. This weighting suppresses detection probability near the coordinate origin—where potential energy is maximal—systematically shifting the detected kinetic energy release (KER) peak away from the statistical mean (Fig. 1, Row 1).
We demonstrate that this masking is governed by an invariant scaling law
$\alpha=\frac{MQ}{\zeta},$ (1)
which couples the fragment mass ($M$), orbital localization ($\zeta$), and repulsive charge ($Q$). We prove that for any localized 3D wavefunction, this coupling constrains the energy detection ratio $R_{E}=E_{\text{peak}}/\langle E\rangle$. Specifically, we identify a 0.33 geometric landmark (Fig. 1, Row 2)—a universal attractor where the detected peak represents approximately one-third of the total energy budget. Our framework predicts a “detection crisis” in highly localized systems: as $\zeta$ increases, $R_{E}$ decays, rendering the true interaction energy effectively invisible to peak-centric calorimetry.
The predictive utility of this scaling law is evidenced by its alignment with diverse experimental benchmarks (Fig. 1, Row 3). In attosecond science, our model resolves the long-standing 7 eV discrepancy in $H_{2}^{+}$ fragmentation, yielding a detection ratio of $R_{E}\approx 0.32$ that matches modern experimental data Zhu (2020, 2021) while accounting for nodal spectral broadening. Remarkably, this molecular landmark is mirrored in the subatomic regime; our distribution provides a robust fit for the historical 210Bi $\beta$-decay spectrum, where the average electron energy remains tethered to approximately one-third of the endpoint energy Ellis and Wooster (1927). This suggests that the 0.33 ratio functions as a universal geometric resonance gate, where energy transfer is dominated by the maximum spatial overlap of 3D wavefunctions. The persistence of this ratio across twelve orders of magnitude identifies the 3D volume element as a universal arbiter of energy detection. Ultimately, this framework establishes a unified baseline for energy reconstruction, providing a rigorous geometric filter to distinguish intrinsic topological masking from the signatures of new physics.
II Theoretical Framework and Numerical Validation
II.1 Sudden Approximation and Bohmian Local Energy
@@
The core of our physical model rests on the sudden approximation, which is valid when the system’s Hamiltonian evolves on a timescale $\tau_{\text{pert}}\ll\hbar/\Delta E$. Under this condition, the initial wavefunction $\psi_{0}(\mathbf{r})$ remains “frozen” at $t=0$, as the probability density has insufficient time to reorganize Makarov (2016). We model the fragmentation event as an instantaneous transition from an attractive bound potential to a repulsive Coulomb configuration:
$V_{\text{initial}}(r)=-\frac{Z_{\text{att}}}{r}\longrightarrow V_{\text{final}}(r)=+\frac{Q}{r}.$ (2)
This framework captures the essential non-adiabatic dynamics of ultrafast processes across scales, from attosecond molecular dissociation to the rapid loss of confinement in nuclear physics (see Sec. A.1 for a detailed comment).
@@
To map the initial-state geometry to final energy observables, we utilize the de Broglie-Bohm formulation of quantum mechanics. In this representation, each point in the initial density corresponds to a trajectory with a conserved total energy $E(\mathbf{r})$, defined as the sum of the external repulsive potential and the local (quantum) kinetic energy density $T_{\text{loc}}(\mathbf{r})=-\frac{1}{2M}\frac{\nabla^{2}\psi_{0}(\mathbf{r})}{\psi_{0}(\mathbf{r})}$. For a generalized Slater-type orbital $R(r)=Nr^{n-1}e^{-\zeta r}$, the analytical expression for the conserved total energy is:
$E(r)=-\frac{1}{2M}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)+\frac{Q}{r}.$ (3)
@@
Due to energy conservation, a particle released from position $r$ retains the energy of its initial state. Consequently, the observed fragment energy distribution $P(E)$ is equivalent to that of the initial state, constructed by sampling these local energies according to the 3D radial probability density $\rho(r)=4\pi r^{2}|\psi_{0}(r)|^{2}$. Crucially, the $r^{2}$ volume weighting acts as a topological filter. In highly localized systems (high $\zeta$), the probability density is concentrated at small $r$, where the local kinetic energy is significantly negative due to high wave-function curvature. This geometric weighting ensures that the mode of the distribution ($E_{\text{peak}}$) is systematically shifted to energies lower than the integrated mean $\langle E\rangle$.
II.2 Comparison with TDSE Benchmarks
@@
To evaluate the predictive power of this 3D analytical model, we benchmarked the resulting $P(E)$ distributions against full-dimensional quantum simulations using the tRecX Time-Dependent Schrödinger Equation (TDSE) solver Scrinzi (2022). While comprehensive comparisons across different grid parameters are provided in Sec. A.3, we emphasize that our analytical local-energy approach captures the primary experimental signatures—specifically the peak energy $E_{\text{peak}}$—with high precision across both the perturbative and saturation regimes. The consistency between the Bohmian peak predictions and the full TDSE results (see Fig. 5(a)) confirms that the $4\pi r^{2}$ volume element, rather than complex dynamical interference, is the dominant driver of the observed energy morphology in 3D Coulomb explosions.
II.3 Parametric Sensitivity and Spectral Morphology
@@
To characterize the influence of physical parameters on the detected energy landscape, we systematically varied the repulsive Coulomb charge $Q$ and the orbital exponent $\zeta$ (Fig. 2). As illustrated in Fig. 2(a), increasing the repulsive charge $Q$ induces a linear translation of the distribution $P(E)$ toward higher energies. This shift is accompanied by a moderate spectral broadening, as a larger $Q$ amplifies the potential gradient across the spatial support of the initial wavefunction.
In contrast, the evolution of $P(E)$ with respect to the orbital exponent $\zeta$ reveals a non-linear sensitivity (Fig. 2(b)). As $\zeta$ increases, the wavefunction contracts, simultaneously elevating the local kinetic energy density—governed by the quadratic scaling $T_{\text{local}}\propto\zeta^{2}$—and shifting the density into regions of steeper potential energy. Notably, the distribution broadens significantly more rapidly with $\zeta$ than with $Q$. This suggests that the internal “quantum pressure” of the initial state is the primary driver of spectral dispersion, outweighing the influence of the external repulsive field. These morphological trends confirm that fragmentation spectra fundamentally function as a probe of the initial-state radial geometry rather than a direct measurement of the integrated energy mean.
II.4 The Energy Detection Ratio $R_{E}$ and the Detection Crisis
To quantify the discrepancy between experimentally observable fragment signals and the total energy budget, we define the dimensionless detection ratio:
@@ II.4 The Energy Detection Ratio $R_{E}$ and the Detection Crisis
where $E_{\text{peak}}$ corresponds to the mode of $P(E)$ and $\langle E\rangle$ represents the integrated expectation value of the total energy. This ratio serves as a diagnostic for the efficiency of peak-centric calorimetry in capturing the true system energy. Figures 2(c) and 2(d) illustrate the systematic evolution of $R_{E}$ across the parametric landscape.
II.4.1 Asymptotic Convergence and Classicalization
As shown by the black curve in Fig. 2(c), $R_{E}$ increases monotonically with the repulsive charge $Q$, asymptotically approaching a universal limit of $0.5$. In this high-$Q$ regime, the repulsive potential $V=Q/r$ dominates the local kinetic energy terms, “classicalizing” the energy landscape and shifting the spectral peak closer to the statistical mean. However, the fact that $R_{E}$ remains strictly bounded by $0.5$ underscores that 3D quantum spatial spread imposes an inescapable topological constraint on the fragment energy spectrum.
II.4.2 The Shell-Dependent Detection Crisis
The decay in $R_{E}$ is observed as the orbital exponent $\zeta$ increases (red curve in Fig. 2(c)). Within our framework, $\zeta$ acts as a proxy for localization and effective nuclear charge; small $\zeta$ values characterize diffuse valence orbitals, while large $\zeta$ represent highly localized core-shell states. As $\zeta$ increases, the quadratic scaling of the quantum pressure near the coordinate origin drives $\langle E\rangle$ to values vastly exceeding $E_{\text{peak}}$, which remains constrained by the most probable radial position.
Our findings identify a shell-dependent detection crisis: while valence-shell fragmentation (low $\zeta$) remains relatively transparent to calorimetry, inner-shell events suffer from massive geometric masking. For $\zeta\approx 5$, a detector may resolve as little as 15% of the true interaction energy. Given that localization typically scales with fragment mass, the extreme confinement of nucleons in nuclear physics implies a significantly suppressed $R_{E}$. The global contour map in Fig. 2(d) synthesizes these trends, confirming that the deeper a state is localized within the potential well, the more effectively 3D quantum geometry masks the true energy budget from classical detection.
III The Dimensionless Master Curve and Scaling Symmetry
@@
The emergence of the 0.33 landmark is rooted in a fundamental scaling symmetry inherent to three-dimensional fragmentation dynamics. To characterize this universality, we examine the peak-to-mean ratio $R_{E}$ across the multi-dimensional parameter space of mass ($M$), orbital localization ($\zeta$), and repulsive charge ($Q$). As demonstrated in Fig. 3, the energy morphology is governed by a single dimensionless coupling constant $\alpha$, defined as:
@@
For a $1s$ Slater-type orbital, substituting this scaling relation reveals that the local kinetic density scales as $\rho_{T}\propto\zeta^{2}/M=\alpha^{-1}\zeta Q$, while the local potential density follows $\rho_{V}\propto Q\zeta$ (see derivation in Sec. A.5). By transforming to a dimensionless radial coordinate $r_{s}=\zeta r$, we find that both energy components at every coordinate point scale linearly with the product $\zeta Q$ for any fixed $\alpha$. This local synchronization ensures that the geometric morphology of the resulting $P(E)$ distribution is invariant to the absolute energy scale. This symmetry is confirmed in Fig. 3(a), where normalized distributions $P(E/Q\zeta)$ for disparate $Q$ and $\zeta$ collapse onto identical master curves. With the $Q\zeta$ scaling of $E$, the distribution converges toward a fixed peak position, providing a robust, scale-invariant baseline for extracting intrinsic physical parameters from fragmented experimental spectra.
Furthermore, this symmetry implies that the detection ratio $R_{E}$ remains invariant along $M\propto\zeta$ trajectories. The system undergoes a self-similar expansion in energy space where the $4\pi r^{2}$ masking effect is preserved as a constant geometric partition of the total energy budget. The invariance observed in the parallel trajectories of Fig. 3(b) and the invariant planes of Fig. 3(c) is a direct consequence of this homogeneous scaling. At the $\alpha\approx 1$ equilibrium, this symmetry manifests as a 1D plateau at $R_{E}\approx 0.33$, identifying the landmark as a structural constant of 3D quantum geometry rather than a transient dynamical variable.
@@
As shown in the master curve in Fig. 3(d), the numerical behavior of $R_{E}$ is characterized by the following asymptotic limits:
@@
At $\alpha=1$, the system achieves a geometric equilibrium where the $4\pi r^{2}$ volume element sequesters exactly two-thirds of the total energy in the high-momentum radial tail, yielding the 0.33 landmark. As $\alpha$ increases (representing high repulsion $Q$ or low mass $M$), the ratio monotonically approaches the $0.5$ universal bound—a regime where the repulsive field dominates the initial-state quantum pressure.
@@
Conversely, as the system moves toward high localization $\zeta$ (decreasing $\alpha$), $R_{E}$ decays precipitously, falling below $0.2$ for $\zeta>5$. This signals a profound detection crisis: the vast majority of interaction energy becomes experimentally invisible to peak-centric calorimetry. This confirms that for sufficiently localized states—such as core-shell electrons or compressed nucleons—the true interaction energy is effectively masked by the topological constraints of 3D space. Consequently, the mapping $R_{E}(\alpha)$ serves as a universal master curve for energy reconstruction across all scales of three-dimensional quantum fragmentation.
IV Discussion: From Molecular Calorimetry to Subatomic Decay